The Watson Foundation, a private not-for-profit entity, starts 2017 with cash of $121,000; contributions receivable (net) of $221,000; investments of $321,000; and land, buildings, and equipment of $221,000. In addition, its unrestricted net assets were $442,000, temporarily restricted net assets were $121,000, and permanently restricted net assets were $321,000. Of the temporarily restricted net assets, 50 percent must be used to help pay for a new building; the remainder is restricted for salaries. No implied time restriction was designated for the building when purchased. For the permanently restricted net assets, all income is unrestricted.

During the current year, the entity has the following transactions:

  • Computed interest of $41,000 on the contributions receivable.
  • Received cash of $121,000 on the contributions and wrote off another $6,100 as uncollectible.
  • Received unrestricted cash gifts of $201,000.
  • Paid salaries of $111,000 with $36,000 of that amount coming from restricted funds.
  • Received a cash gift of $33,000 that the entity must convey to another charity. However, Watson has the right to give the money to a different organization if it so chooses.
  • Bought a building for $521,000 by signing a long-term note for $460,500 and using restricted funds for the remainder.
  • Collected membership dues of $51,000. Individuals receive substantial benefits from the memberships.
  • Received income of $51,000 generated by the permanently restricted net assets.
  • Paid rent of $33,000, advertising of $36,000, and utilities of $37,000.
  • Received an unrestricted pledge of $221,000; it will be collected in five years. The organization expects to collect the entire amount. Present value is $170,000. It then recognized interest of $8,100 for the year.
  • Computed depreciation as $61,000.
  • Paid $36,000 in interest on the note signed to acquire the building.
  1. Prepare a statement of activities for this entity for this year.

  2. Prepare a statement of financial position for this entity at the end of this year.
Answer #1

Answer a

Particulars Unrestricted Net Assets Temporarily Restricted Net Assets Permanently Restricted Net Assets
Support
Contributions 201,000 203,000 (170,000+33,000)
Revenue
Membership dues 51,000
Investment income 51,000
Interest income 41,000 8,100
Total Support and Revenue 344,000 211100
Net assets released from restriction 96500(36000+(521000-460500)) (96500)
Total revenues and net assets released from restriction 440,500 114,600
Expenses
General and administrative
Rent 33,000
Salary 111,000
Advertising 36,000
utilities 37,000
Depreciation 61,000
Interest 36,000
Total expenses 314,000
Excess of total revenues and net assets released from restriction over expenses 126,500 114,600
Net assets at beginning of year 442,000 121,000 321,000
Net assets at end of year 568,500 235,600 321,000

STATEMENT OF FINANCIAL POSITION

Amount $ Total Amount $
ASSETS
Cash 264,500
Pledges receivable (net) 319,100
Investments 321,000
Land, buildings, and equipment 742,000
Depreciation 61,000 681000
Total assets 1585600
LIABILITIES
Notes Payable 460500
NET ASSETS
Unrestricted 568,500
Temporarily Restricted 235,600
Permanently Restricted 321,000 1,125,100
TOTAL LIABILITIES AND NET ASSETS 1585600

Cash = 121000+121000+201000-111000+33,000-(521000-460500)+51000+51000-33000-36000-37000-36000 = 264500

Pledges receivable = 221000+41000+170000+8100 = 440100-121000=319,100

Land and buildings = 221000+521000=742000